4. A helicopter drops a supply package to soldiers in a jungle clearing. When the package is dropped, the helicopter is 100 m above the clearing and flying at 35 m/s at an angle of 33° above the horizontal. (a) Where does the package land? (b) If the helicopter flies at constant velocity, where is it when the package lands?
Answer #1

here,

the speed of helicopter , v = 35 m/s

theta = 33 degree

the height of helicopter , h = 100 m

a)

let the time taken to reach the ground be t

for vertical direction ,

- h = v * sin(theta) * t - 0.5 * g * t^2

- 100 = 35 * sin(33) * t - 0.5 * 9.81 * t^2

solving for t

t = 6.86 s

the horizontal distance travelled before reaching the ground , x = v * cos(theta) * t

x = 35 * cos(33) * 6.86 m

x = 201.4 m